# Harm Amplification in Text-to-Image Models

### v2

@@ Introduction
The Distribution-Based Thresholds method utilizes (1) a text safety classifier that is applied to the input prompt to produce text harm scores (the harm severity for a given prompt), and (2) an image safety classifier that is applied on the output image to generate image harm scores (the harm severity for a given image). Consider a set of text harm scores ($T$) and image harm scores ($I$). We first discretize $T$ into $n$ even buckets ($B_{0},B_{1},\dots,B_{n}$) where $l_{j}$ is the lower bound of the bucket and $l_{j+1}$ is the upper bound of the bucket.
$B_{j}=\{t\in T:l_{j}<t\leq l_{j+1}\},\quad j=0,1,...,n$ (1)
For each $B_{j}$, the corresponding image harm scores ($I_{j}$) form a distribution ($D_{j}$). Statistical measures, including mean ($\mu_{j}$) and standard deviation ($\sigma_{j}$), are calculated for each $D_{j}$. The non-fitted, raw image harm amplification threshold for each bucket is determined as $\mu_{j}+2\cdot\sigma_{j}$, or the 95th percentile of the distribution. To provide a smooth representation of the calculated thresholds, a polynomial function can be fitted through the raw threshold values across all buckets. In our experiments, we fit a first-degree polynomial with the corresponding equation to determine the image harm amplification threshold $HarmAmpThresh_{j}$ for each text harm bucket $B_{j}$:
$HarmAmpThresh_{j}=\hat{y}_{j}=b_{1}j+b_{0},$ (2)
$where\quad y_{j}=\mu_{j}+2\cdot\sigma_{j}$
For example, consider a T2I prompt-image dataset where we have image harm scores $I$ and text harm scores $T$. Say we now have a new prompt-image pair where we would like to determine whether harm amplification occurred. Given an image harm score (i.e., $i=0.5$) and text harm score (i.e., $t=0.15$), we categorize the text harm score into a specific bucket (i.e., $B_{1}=0.1<t\leq 0.2$). For that given bucket, the image harm amplification threshold (i.e., $HarmAmpThresh_{1}$) is obtained by fitting a 1-degree polynomial to the 95th percentile threshold for a prior image harm distribution in each bucket. We then compare the corresponding image harm score with the harm amplification threshold to determine whether harm is amplified (i.e., $i>HarmAmpThresh_{1}$).
As mentioned previously, this method works best when there is access to large amounts of data to calculate statistical measures from. Additionally, this method performs well when text and image safety classifier are not well calibrated (i.e., the output harm severity for an image harm classifier does not equate to the severity of harm outputted by the text harm classifier). Rather than comparing text and image harm scores directly which requires well-aligned classifiers, we determine whether an image amplifies harm by comparing its harm score to the relevant distribution within its corresponding text harm bucket, avoiding direct text-image harm score comparisons.
@@
Utilizing text and image harm safety scores similar to Method 1, the Bucket Flip method provides a more direct comparison of discretized harm categories across image and text. Consider a set of text harm scores ($T$) and image harm scores ($I$), assuming text and image harm scores have a similar scale. In cases where the scales differ, techniques such as z-score transformations can be applied to align the scales. We can then discretize $T$ and $I$ into the same $n$ even buckets ($B_{0},B_{1},\dots,B_{n}$) where $l_{j}$ is the lower bound and $l_{j+1}$ is the upper bound of the bucket (see Equation 1).
For a given (prompt, image) pair $(t,i)$, we categorize the text harm score into a specific bucket $B_{t}$ and the image harm score into a specific bucket $B_{i}$. If the image harm bucket is higher than the text harm bucket ($B_{i}>B_{t}$), then we determine that harm amplification occurred. For example, suppose we have discretized harm scores into 10 buckets of ($0-0.1,0.1-0.2,\text{…},0.9-1.0$). Given an image harm score (i.e., $i=0.45$) and text harm score (i.e., $t=0.15$), the image score would fall into 5th bucket and the text score is in the 2nd bucket. As the image score bucket is higher than text score bucket, we determine that there was a bucket flip and harm amplification has occurred.
Unlike Method 1, this method does not require any data and is most effective when both text and image harm classifiers scores are well aligned or when users do not have access to raw harm scores. In many real-world scenarios, safety classifiers often output discretized categories (i.e., low harm, medium harm, high harm) rather than raw harm scores, making the Bucket Flip method a practical approach for identifying harm amplification.
@@
While training dedicated image or text harm classifiers enables the most direct harm amplification measurement methods, many researchers and practitioners may not have the resources to label such data and thus may not be able to train the necessary safety classifiers. In this case, existing large pre-trained image-text models can serve as a substitute. Large pre-trained image-text models, such as CLIP (Radford et al. 2021), embed images and texts such that the cosine distance between corresponding text and images is minimized in the embedding space. Taking advantage of the co-embedding space, harm amplification can be measured by how much closer a generated image’s embedding, $i$, is to a harmful concept in the embedding space than the prompt’s embedding, $p$, is. Although we cannot determine the precise embedding of harm concepts, we can embed words related to these harm concepts (e.g., for the sexually explicit harm category, we can use the words “sexual”, “porn”, etc. - see Appendix, Table 2), defining their text embeddings as $h_{0},h_{1},h_{2},...h_{k}$. These harm word embeddings are meant to approximate where harm concepts such as sexually explicit may lie in the CLIP embedding space. We can then calculate the average distance between the proximity of image embeddings and harm concepts compared to prompt embeddings and harm concepts:
$D_{harm}=\mu(cos\_similarity(h_{k},i)\\ -cos\_similarity(h_{k},p))\mid k\in\{0,1,2,...,K\}$ (3)
@@
Harm amplification is determined to occur if $D_{harm}$ is greater than some predefined harm amplification threshold. This conceptually maps to the notion that harm amplification can be quantified by examining how much closer the image is to the approximated harm concepts compared to the prompt.
Experiments
@@
In this section, we demonstrate how to apply our framework empirically using a combination of safety classifiers, embedding models, and measurement dataset. These methods are then evaluated on an independent human annotated dataset where f1-scores, precision, and recall are derived.
Measurement Dataset and Models
To derive statistics for Method 1, we obtained a measurement dataset consisting of 497,157 prompts representing various demographics and prompt categories. These prompts were sampled from dogfood user data aimed at testing generative AI models. By employing this diverse and representative prompt dataset over curated adversarial or safety-specific prompts, we prioritize a broad spectrum of potential harmful representations that may be encountered in actual deployment. While curated adversarial prompts can be valuable for specific safety assessments, our method takes a broader perspective by aiming to understand how T2I models can generate harm amplifying images over a large range of prompts (from benign to borderline to harmful). We used Stable Diffusion 2.1 (Rombach et al. 2022) to generate images for our prompt measurement dataset. Four images were generated per prompt resulting in 1,988,628 total generated images. No NSFW or safety filters were applied during the image generation process. Thus, potentially harmful images were generated for our experiment and were not filtered or blocked.
Safety machine learning classifiers were used to provide harm scores for sexually explicit content and violent content for Methods 1 and 2. Separate classifiers were used on the input text prompts and the output images with harm scores for each classifier ranging from 0 to 1. For the pre-trained image-text model used in Method 3, we chose CLIP Resnet 101 (Radford et al. 2021). In addition to safety classfiers and image-text embedding models, fairness machine learning classifiers (Baruah et al. 2022) that identified probability estimates of gender presentation in images were also used with estimates for each gender (male, female, or undefined) ranging from 0 to 1 in the evaluation data.
@@
To create the evaluation dataset, we sourced 1125 prompt, image pairs from the Adversarial Nibbler data challenge (Quaye et al. 2024). This public, crowd-sourced challenge aimed to identify failure-modes of DALL-E 2 (Open AI 2023) and various Stable Diffusion versions (XL 1.0, 1.5, and 2.1) (Stability AI 2022). Five to six raters trained on AI safety policies annotated prompts and images separately for the presence of sexually explicit content or violent content. To derive a confidence score for sexually explicit and violence, we converted the ratings of the annotators into a proportional measure (i.e., if 3 out of 5 raters rated an image as violent, the violence score for that image would be 0.6 or 60%). The ground truth labels for harm amplification for each harm were then defined as whether the image confidence score for a given harm was greater than the corresponding text confidence score (i.e., if 80% of raters rated the image violent and only 60% rated the corresponding prompt violent, then the image amplified harm). If the image confidence score was less than or equal to the corresponding text confidence score for a given harm, then the ground truth was labeled as harm amplification not occurring. We calculated precision, recall, and f1-scores for each method’s evaluation of harm amplification on the Nibbler dataset. This evaluation was conducted separately for sexually explicit and violence harm amplification.
Furthermore, we conducted an additional analysis on the evaluation dataset to examine gender fairness. Specifically, we investigated whether harm amplification of sexually explicit or violent content disproportionately affected one perceived gender. Perceived gender was machine annotated in the evaluation dataset with significance testing performed on rates of harm amplification across genders.
Experimental Results
Evaluation of Methods in Detecting Harm Amp
To quantify harm amplification using Method 1, Distribution-Based Threshold Method, we discretized the text harm scores for the measurement prompt data into 5 buckets for each harm type: sexually explicit and violence. Subsequently, we then obtained an image harm distribution from the measurement data for each of the 5 buckets corresponding to each harm type (see Appendix, Figure 3A). The 95th percentile score, or 2 standard deviations above the mean, was then calculated for each distribution within its respective bucket. A first-degree polynomial was fitted across buckets on the 95th percentile score resulting in new fitted thresholds (Appendix, Figure 3B) for each harm type. For our data, a first-degree linear function was most appropriate, but the type of function fitted on thresholds may vary depending on the statistics of the image data distribution. Having derived fitted image harm amplification thresholds for each text bucket for each harm type, we then assessed how well our measurement technique performed on the Nibbler evaluation dataset. A precision of 0.864 and recall of 0.985 was observed when evaluating sexually explicit content in the Nibbler dataset (see Table 1). When evaluating violent harm amplification, we noted a precision of 0.460 and recall of 0.831.
@@
While we expect Method 3, Image-Text Co-embedding Method, to perform worse at the task of identifying instances of harm amplification than using dedicated sexually explicit and violence text and image classifiers, this method is useful for those with access to limited resources. We used CLIP to encode images and prompts, and defined sexually explicit harm concept using 15 words such as “porn”, “sexual”, and “nude” (see Appendix, Table 2). These concept words were sourced from internal adversarial testing and (Rando et al. 2022a) in their attempt to reverse-engineering Stable Diffusion’s NSFW filter. Adversarial testing was used to find 15 violence concept words such as “violence”, “weapons”, and “blood” (see Appendix, Table 2). Various harm amplification thresholds were applied to the evaluation dataset to obtain precision-recall curves (see Appendix, Figure 5). Unlike the other methods, there is not an obvious way to define what the best harm amplification threshold is, but rather, practitioners can decide on an appropriate threshold based on their specific use case. For these results we obtained a threshold that maximized the f1 score, and computed precision and recall from this number using the PR curves. As with other methods, we saw better results for sexually explicit harm amplification (precision of 0.585, recall of 0.920) than violence harm amplification (precision 0.322, recall 0.566).
Gender Disparities in Harm Amplification
@@
Machine annotations of perceived gender (Baruah et al. 2022) were collected for each image in the Nibbler evaluation dataset. We restricted our analysis to images in which there was a strict majority of a perceived gender. Images in which there were no faces or where the number of perceived males equaled that of perceved females were discarded. We then calculated the rates of harm amplification using the majority perceived gender. Our analysis reveal that perceived females have significantly higher rates of oversexualization ($p<0.001$, see Figure 2), in line with previous literature positing theories of gendered sexual objectification stereotypes (Fredrickson and Roberts 1997; Szymanski, Moffitt, and Carr 2011). There was no significant difference in violence harm amplification across perceived genders ($p=0.492$). We additionally show how well each method performed in detecting harm amplification across genders in the appendix section (see Table 3).
@@
Limitations
While this framework is shown to be effective in identifying harm amplification, we acknowledge the limitations of our methodologies and identify potential areas of future research. Method 1, the Distribution-Based Thresholds Method is highly sensitive to the measurement dataset’s distribution of scores. If the distribution is skewed (i.e., the dataset is predominantly either safer or more harmful), derived harm amplification thresholds will then be biased. Thus, when using this method, it is imperative to analyze and verify that the distribution of harm scores within the measurement dataset aligns with practitioners’ assumptions. As noted previously, Method 2, the Bucket Flip Method relies on the successful calibration of text and image safety classifiers. Uncalibrated classifiers can result in inaccurate identification of harm amplification. Future research should investigate how calibration methods and various types of safety classifiers affect this method. Method 3, the Image-Text Co-embedding Method, utilized the open source model CLIP to find the difference in embeddings of images and prompts in relation to harm concept words. However, CLIP has been shown to exhibit fairness issues particularly concerning demographic attributes such as gender and race (Luo et al. 2024; Wang, Zhang, and Sang 2022). A potential future direction of work could involve using a debiased version of CLIP similar to that of FairCLIP (Wang, Zhang, and Sang 2022) to investigate its efficacy in identifying harm amplification instances.
Across all methods, we see higher performance in identifying instances of harm amplification in sexually explicit content compared to violent content. Further research should investigate the factors contributing to this disparity in performance such as differences in safety classifier effectiveness, noise in human annotations, or biases within datasets used.

### v3

@@ Methods
The Distribution-Based Thresholds method utilizes: 1) a text safety classifier that is applied to the input prompt to produce text harm scores (the harm severity for a given prompt), and 2) an image safety classifier that is applied on the output image to generate image harm scores (the harm severity for a given image). Consider a set of text harm scores (${H(t):t\in T}$) and image harm scores (${H(i):i\in I}$). We first discretize the text harm scores into $n$ even buckets ($B_{0},B_{1},\dots,B_{n}$) where $l_{j}$ is the lower bound of the bucket and $l_{j+1}$ is the upper bound of the bucket.
$B_{j}=\{t\in T:l_{j}<H(t)\leq l_{j+1}\},\quad j=0,1,...,n-1$ (2)
For each $B_{j}$, the corresponding image harm scores $H(i):i=M(t),t\in B_{j}$ form a distribution $D_{j}$. Statistical measures, including mean ($\mu_{j}$) and standard deviation ($\sigma_{j}$), are calculated for each $D_{j}$. The non-fitted, raw image harm amplification threshold for each bucket is determined as $\mu_{j}+2\cdot\sigma_{j}$, or the 95th percentile, $P_{95,j}$ of the distribution. To provide a smooth representation of the calculated thresholds, a polynomial function can be fit through the raw threshold values across all buckets. In our experiments, we fit a first-degree polynomial with the corresponding equation to determine the image harm amplification threshold $HarmAmpThresh_{j}$ for each text harm bucket $B_{j}$:
$HarmAmpThresh(j)=b_{1}j+b_{0}$ (3)
where $b_{1}$ and $b_{0}$ are the coefficients of the fitted polynomial, determined by minimizing the least squares error:
$\min_{b_{1},b_{0}}\sum_{j=0}^{n-1}(P_{95,j}-(b_{1}j+b_{0}))^{2}$ (4)
Harm amplification is determined to occur for a pair $(t,i)$ if $t\in B_{j}$ and $H(i)>HarmAmpThresh(j)$. As mentioned previously, this method works best when there is access to large amounts of data to calculate statistical measures from. Additionally, this method performs well when text and image safety classifier are not well calibrated (i.e., the output harm severity for an image harm classifier does not equate to the severity of harm outputted by the text harm classifier). Rather than comparing text and image harm scores directly which requires well-aligned classifiers, we determine whether an image amplifies harm by comparing its harm score to the relevant distribution within its corresponding text harm bucket, avoiding direct text-image harm score comparisons.
@@
Utilizing text and image harm safety scores similar to Method 1, the Bucket Flip method provides a more direct comparison of discretized harm categories across image and text. Consider a set of text harm scores (${H(t):t\in T}$) and image harm scores (${H(i):i\in I}$) obtained by safety classifiers. Assuming that the classifiers are well aligned, we can discretize text and image harm scores into the same $n$ even buckets ($B_{0},B_{1},\dots,B_{n}$) where $l_{j}$ is the lower bound and $l_{j+1}$ is the upper bound of the bucket (see Equation 2).
For a given (prompt, image) pair $(t,i)$, we categorize the text harm score into a specific bucket $B_{t}$ and the image harm score into a specific bucket $B_{i}$. If the image harm bucket is higher than the text harm bucket ($B_{i}>B_{t}$), then we determine that harm amplification occurred.
Unlike Method 1, this method does not require any data and is most effective when both text and image harm classifiers scores are well aligned or when users do not have access to raw harm scores. In many real-world scenarios, safety classifiers often output discretized categories (i.e., low harm, medium harm, high harm) rather than raw harm scores, making the Bucket Flip method a practical and easy approach for identifying harm amplification.
@@
While training dedicated image or text harm classifiers enables the most direct harm amplification measurement methods, many researchers and practitioners may not have the resources to label such data and thus may not be able to train the necessary safety classifiers. In this case, existing large pre-trained image-text models can serve as a substitute. Large pre-trained image-text models, such as CLIP (Radford et al. 2021), embed images and their corresponding text labels such that the cosine distance between corresponding text and images is minimized in the embedding space. Taking advantage of the co-embedding space, harm amplification can be measured by how much closer a generated image’s CLIP embedding, $z_{i}$, is to a harmful concept in the embedding space than the prompt’s CLIP embedding, $z_{t}$, is. Although we cannot determine the precise embedding of harm concepts, we can embed words related to these harm concepts (e.g., for the sexually explicit harm category, we can use the words “sexual”, “porn”, etc. - see Appendix, Table 2), defining the texts as $\{h_{k}\}_{k=1}^{K}$ and their subsequent embeddings as $z_{h_{k}}$ These harm word embeddings are meant to approximate where harm concepts such as sexually explicit may lie in the CLIP embedding space. We can define our harm function $H$ by calculating the average cosine similarity between a given text or image embedding and harm concepts:
$H(x)=\frac{1}{K}\sum_{k=1}^{K}cos\_sim(z_{h_{k}},z_{x})$ (5)
where $x$ can denote a text or an image and $cos\_sim(z_{h_{k}},z_{x})=\frac{\mathbf{z}_{h_{k}}\cdot\mathbf{z}_{x}}{\|\mathbf{z}_{h_{k}}\|\|\mathbf{z}_{x}\|}$. Harm amplification is then determined to occur for a pair $(t,i)$ if:
$H(i)-H(t)>\tau$ (6)
@@
where $\tau>0$ is a predefined threshold. This conceptually maps to the notion that harm amplification can be quantified by examining how much closer the image is to the approximated harm concepts compared to the prompt. It is of note, that there may exist biases in the CLIP embedding space that may influence what is considered harm amplification. Applying this method using a debiased embedding space (Wang, Zhang, and Sang 2022) should be considered as a future direction.
Evaluation of Methods
@@
Here, we demonstrate how to apply our framework empirically using a combination of safety classifiers, embedding models, and measurement dataset. These methods are then evaluated on an independent human annotated dataset where f1-scores, precision, and recall are derived.
@@
Method 1 Measurement Dataset
To derive statistics for Method 1, we obtained a measurement dataset consisting of 497,157 prompts representing various demographics and prompt categories. These prompts were sampled from dogfood user data aimed at testing generative AI models. By employing this diverse and representative prompt dataset over curated adversarial or safety-specific prompts, we prioritize a broad spectrum of potential harmful representations that may be encountered in actual deployment. While curated adversarial prompts can be valuable for specific safety assessments, our method aims to understand how harm amplification can occur over a large range of prompts (from benign to borderline to harmful). We used Stable Diffusion 2.1 (Rombach et al. 2022) to generate images for our prompt measurement dataset. Four images were generated per prompt and no NSFW or safety filters were applied during the image generation process. Thus, potentially harmful images were generated for our experiment and were not filtered or blocked.
Safety Classifiers and Image-Text Models
Pretrained safety machine learning classifiers were used to provide harm scores for sexually explicit content and violent content for Methods 1 and 2. Separate classifiers were used on the input text prompts and the output images with harm scores for each classifier ranging from 0 to 1. For the pre-trained image-text model used in Method 3, we chose CLIP Resnet 101 (Radford et al. 2021).
@@
To create the evaluation dataset, we sourced 1125 prompt, image pairs from the Adversarial Nibbler data challenge (Quaye et al. 2024). This public, crowd-sourced challenge aimed to identify failure-modes of DALL-E 2 (Open AI 2023) and various Stable Diffusion versions (XL 1.0, 1.5, and 2.1) (Stability AI 2022). Five to six raters trained on AI safety policies annotated prompts and images separately for the presence of sexually explicit content or violent content. To derive a confidence score for sexually explicit and violence, we converted the ratings of the annotators into a proportional measure (i.e., if 3 out of 5 raters rated an image as violent, the violence score for that image would be 0.6 or 60%). The ground truth labels for harm amplification were then defined as whether the image confidence score for a given harm was greater than the corresponding text confidence score (i.e., if 80% of raters rated the image violent and only 60% rated the corresponding prompt violent, then the image amplified harm). We calculated precision, recall, and f1-scores for each method’s evaluation of harm amplification on the Nibbler dataset. This evaluation was conducted separately for sexually explicit and violence harm amplification.
Evaluation of Methods in Detecting Harm Amplication
To quantify harm amplification using Method 1, Distribution-Based Threshold Method, we discretized the text harm scores for the measurement prompt data into 5 buckets for each harm type: sexually explicit and violence. Subsequently, we then obtained an image harm distribution from the measurement data for each of the 5 buckets corresponding to each harm type (see Appendix, Figure 5A). The 95th percentile score was then calculated for each distribution within its respective bucket with a first-degree polynomial being fit across buckets on the 95th percentile score resulting in new fitted thresholds (Appendix, Figure 5B) for each harm type. For our data, a first-degree linear function was most appropriate, but the type of function fitted on thresholds may vary depending on the statistics of the image data distribution. Having derived fitted image harm amplification thresholds for each text bucket for each harm type, we then assessed how well our measurement technique performed on the Nibbler evaluation dataset. A precision of 0.864 and recall of 0.985 was observed when evaluating sexually explicit content in the Nibbler dataset (see Table 1). When evaluating violent harm amplification, we noted a precision of 0.460 and recall of 0.831.
@@
While we expect Method 3, Image-Text Co-embedding Method, to perform worse at the task of identifying instances of harm amplification than using dedicated sexually explicit and violence text and image classifiers, this method is useful for those with access to limited resources. We used CLIP to encode images and prompts, and defined sexually explicit harm concept using 15 words such as “porn”, “sexual”, and “nude” (see Appendix, Table 2). These concept words were sourced from internal adversarial testing and (Rando et al. 2022a) in their attempt to reverse-engineering Stable Diffusion’s NSFW filter. Adversarial testing was used to find 15 violence concept words such as “violence”, “weapons”, and “blood” (see Appendix, Table 2). Various harm amplification thresholds were applied to the evaluation dataset to obtain precision-recall curves (see Appendix, Figure 7). Unlike the other methods, there is not an obvious way to define what the best harm amplification threshold is, but rather, practitioners can decide on an appropriate threshold based on their specific use case. For these results we obtained a threshold that maximized the f1 score, and computed precision and recall from this number using the PR curves. As with other methods, we saw better results for sexually explicit harm amplification (precision of 0.585, recall of 0.920) than violence harm amplification (precision 0.322, recall 0.566). Poorer performance on violence across methods could be attributed to less accurate classification models or noisier human labels. Future research should explore biases within classifiers and datasets and their impact on measuring harm amplification.
Examining Patterns of Harm Amplification
In the previous section, we showed how to apply different methodologies to measure harm amplification. Here, we explore patterns of harm amplification within the data, examining gender differences and applying explainability techniques to gain deeper insights of how different inputs affect harm amplification.
Gender Disparities
@@
Annotations of perceived gender were collected for each image in the Nibbler evaluation dataset using gender classifiers (Baruah et al. 2022). We restricted our analysis to images in which there was a strict majority of a perceived gender. Images in which there were no faces or where the number of perceived males equaled that of perceved females were discarded. We then calculated the rates of harm amplification using the majority perceived gender. Our analysis reveal that perceived females have significantly higher rates of oversexualization ($p<0.001$, see Figure 2),
@@
in line with previous literature positing theories of gendered sexual objectification stereotypes (Fredrickson and Roberts 1997; Szymanski, Moffitt, and Carr 2011). There was no significant difference in violence harm amplification across perceived genders ($p=0.492$). We additionally show how well each method performed in detecting harm amplification across genders in the appendix section (see Table 3).
Applying Explainability Techniques
To gain deeper insights into the mechanisms of harm amplification, we employed explainability techniques such as attribution maps and counterfactual analysis. We utilized Diffusion Attentive Attribution Maps (DAAM) (Tang et al. 2023) to visualize the relationship between input words and output image pixels, leveraging cross attention maps in diffusion models. Figure 3 shows attribution maps for two examples in our dataset that exhibited notable harm amplification.
Figure 3: Attribution maps (Tang et al. 2023) for notable harm amplification prompts.
For the prompt “Person venting frustration.”, the attribution maps show that the words “frustration” and to a lesser extent “venting” contribute to the blood splatter observed in the image. In contrast, for the prompt “black gay man”, while the word “gay” seems to be attributed to some facial and neck features, the attribution maps alone do not capture what aspects of the prompt contribute the man being depicted shirtless.
To further investigate cases where attribution maps provide limited insights, we ran a counterfactual analysis. For the prompt “black gay man”, we generated counterfactual variations by altering race and sexual orientation terms. Using the same generation seed, we produced images from these counterfactual prompts (Figure 4).
Our analysis revealed that prompts altering the race term still exhibited similar levels of harm amplification. However, when modifying the sexual orientation term, we observed a significant reduction in harm amplification suggesting that the sexual orientation term played a more substantial role in triggering sexually explicit harm amplification than the race term. These findings show the complex interaction between different demographic descriptors and harm amplification in T2I models and highlight the need for future research of intersectional identities and harm amplification.
Figure 4: Counterfactual analysis for the prompt “black gay man”, altering the sexual orientation term (red) and the race term (blue).